Calculate the Sodium carbonate

content of 1.12g sample requiring

24.8ml of 0.1947N HCL on titration

with phenolphthalein


1
Expert's answer
2021-02-23T04:31:05-0500

NaHCO3 + HCl = NaCl + H2O + CO2

n (NaHCO3) = n (HCl)

Cn = N/V

N (HCl) = N (NaHCO3) = V x Cn = 0.0248 x 0.1947 = 0.005 g-equivalents


Number of gram equivalents = weight of solute × [Equivalent weight of solute]-1

Equivalent weight of solute = M/number of equivalents

M (NaHCO3) = 84.007 g/mol

Number of equivalents (NaHCO3) = 2

Equivalent weight of solute (NaHCO3) = 84.007/2 = 42.007


Weight of solute (NaHCO3) = Number of gram equivalents x Equivalent weight of solute = 0.005 x 42.007 = 0.21 g


%(NaHCO3) = 0.21/1.12 x 100 = 18.75%